### v1

# Secure Smart Contract with Control Flow Integrity

Abstract.
Smart contracts power decentralized financial (DeFi) services but are vulnerable to complex security exploits that can lead to significant financial losses. Existing security measures often fail to adequately protect these contracts due to the composability of DeFi protocols and the increasing sophistication of attacks. Through a large-scale empirical study of historical transactions from the $30$ hacked DeFi protocols, we discovered that while benign transactions typically exhibit a limited number of unique control flows, in stark contrast, attack transactions consistently introduce novel, previously unobserved control flows. Building on these insights, we developed CrossGuard, a novel framework that enforces control flow integrity in real-time to secure smart contracts. Crucially, CrossGuard does not require prior knowledge of specific hacks; instead, it dynamically enforces control flow whitelisting policies and applies simplification heuristics at runtime. This approach monitors and prevents potential attacks by reverting all transactions that do not adhere to the established control flow whitelisting rules. Our evaluation demonstrates that CrossGuard effectively blocks $28$ of the $30$ analyzed attacks when configured only once prior to contract deployment, maintaining a low false positive rate of $0.28\%$ and minimal additional gas costs. These results underscore the efficacy of applying control flow integrity to smart contracts, significantly enhancing security beyond traditional methods and addressing the evolving threat landscape in the DeFi ecosystem.
copyright: nonecopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
1. Introduction
Blockchain technology has revolutionized the creation of global, secure, and programmable ledgers, fundamentally altering how digital transactions are conducted. Central to this innovation are smart contracts, which operate on blockchains, allowing developers to define and enforce complex transactional rules directly on the ledger. This capability has positioned blockchains and smart contracts as the backbone of various decentralized financial (DeFi) services. As of March 13th, 2025, the total value locked in 3,973 DeFi protocols has surged to approximately \$87.82 billion (DefiLlama, 2025a), highlighting the substantial economic impact and growth of this technology.
However, the increasing reliance on smart contracts has brought security concerns to the forefront. By the same date, vulnerabilities in DeFi smart contracts have resulted in financial losses exceeding \$11.21 billion USD (DefiLlama, 2025b). In response to these growing threats, researchers have developed an array of program analysis and verification techniques and tools aimed at securing smart contracts (Luu et al., 2016; Torres et al., 2018; Liu et al., 2018; Feist et al., 2019; Zhang et al., 2019; Bose et al., 2022; Ma et al., 2021; Zheng et al., 2022; Wang et al., 2024; Albert et al., 2020; Zhang et al., 2023). Additionally, to mitigate risks and ensure the integrity of smart contracts, developers often commission thorough security audits to identify and address potential errors and vulnerabilities prior to the contract deployment.
Despite these advancements in security measures, the evolving landscape of smart contracts has continually outpaced traditional defenses. Modern smart contracts are now designed with the flexibility to support the layering of additional contracts, a feature particularly vital in the decentralized finance (DeFi) ecosystem. In DeFi, smart contracts facilitate a diverse array of financial products and services, including lending, borrowing, trading, and yield farming. These interlinked contracts are often referred to as “DeFi legos,” emphasizing their modularity. The ability to combine various DeFi smart contracts, a concept known as “DeFi composability,” is widely regarded as one of the key advantages of DeFi (Popescu, 2020; Amler et al., 2021; Schär, 2021). However, this complexity and interdependence introduce significant challenges to securing smart contracts with conventional methods. The security of a DeFi protocol depends not only on the correct design and implementation of its own contracts but also on the integrity of external contracts it interacts with. Additionally, experienced users or attackers can deploy their own smart contracts to invoke functions across multiple DeFi protocols in arbitrary sequences. Considering all possible interactions a DeFi protocol may encounter prior to deployment is often infeasible for traditional security approaches.
A critical observation in hack transaction analysis is that they often exploit unintended control flows across multiple functions, deviating from the original design intentions of the developers. For instance, re-entrancy attacks leverage an unforeseen recursive control flow through default handlers in custom contracts, enabling repeated execution of a critical function within one transaction. Similarly, flash loan attacks manipulate multiple functions in a precisely timed sequence, utilizing large asset transfers to coerce the victim contract into executing unfavorable trades.
To further explore this phenomenon, we conducted an empirical study analyzing in total $1,327,925$ historical transactions of $30$ compromised protocols on Ethereum. Our findings reveal that the number of different control flows is relatively constrained. In all but two cases, attack transactions introduced novel control flows that had never been observed in any previous transaction.
CrossGuard: Building on the above observations, we developed CrossGuard, a novel framework to enforce control flow integrity to secure smart contracts. Given a DeFi protocol, CrossGuard instruments its existing smart contracts with additional code to track control flow data. CrossGuard also deploys a new guard contract that collects control flow data from these instrumented contracts, and enforces four whitelisting policies at runtime to detect and neutralize any attacks that attempt to exploit unexpected control flows. Unlike many previous invariant enforcement tools (Chen et al., 2024; Liu and Li, 2022), CrossGuard does not rely on inferring its security rules from prior benign transaction traces and therefore can apply to smart contracts immediately at their initial deployments, leaving no gap of unprotected periods.
Note that control flow integrity is a well-established security technique in traditional software to prevent memory attacks from hijacking a program’s control flow (Burow et al., 2017; Abadi et al., 2009; Göktas et al., 2014). Though our proposed techniques in this paper are dramatically different from prior control flow integrity techniques for traditional software, we adapt this concept here to emphasize the critical role of control flow in our approach.
A key challenge for CrossGuard arises from its fundamental design to only whitelist control flows rather than using prior knowledge of hacks to blacklist potentially malicious ones. This approach, while enhancing security by adhering strictly to known safe paths, could inherently lead to an increase in false positives if not meticulously managed. Although the number of unique control flows is inherently limited, a naive whitelisting approach could still produce numerous false positives or demand substantial human intervention. To address this issue, CrossGuard employs a set of heuristics designed to simplify the collected control flows. Specifically, CrossGuard excludes all read-only function calls from the control flows, as these calls do not alter the blockchain ledger’s state. Additionally, CrossGuard records the ledger state accessed by each function call and tracks the read-after-write dependencies of these calls. If function calls in a control flow trace do not have any dependencies with each other, CrossGuard will treat them as separate control flows and assess them with whitelisting policies individually. These heuristics effectively reduce redundant traces and significantly lower the false positive rate for CrossGuard.
Experimental Results: We evaluated CrossGuard on the deployed smart contracts and their transactions of the $30$ DeFi protocols included in our empirical study. Our results indicate that, when configured only once before deployment, CrossGuard can effectively prevent $28$ out of $30$ attacks analyzed in our study, maintaining a low average false positive rate of just 0.28%. Unlike traditional methods, CrossGuard does not depend on historical transactions. Despite this, CrossGuard still surpasses the state-of-the-art which instruments the smart contracts with invariants learned from historical transactions. Moreover, after implementing two optimization techniques, CrossGuard achieves a minimal gas consumption overhead of 15.52% on average. These results demonstrate the usefulness of our empirical findings and CrossGuard.
Contributions: This paper makes the following contributions.
  • Empirical Study: To the best of our knowledge, we conducted the first comprehensive empirical study of control flows in historical transactions of compromised DeFi protocols. Our analysis uncovers critical insights into the control flow patterns prevalent in DeFi protocols and explores various use cases of DeFi composability.
  • CrossGuard Technique: This paper proposes the first control flow integrity technique for smart contracts with whitelisting policies and simplification heuristics. This paper also details methods for implementing these policies and heuristics through static and dynamic analysis, and describes how they are instrumented in contracts and enforced on the fly.
  • Evaluation and Tools: This paper evaluates the effectiveness of CrossGuard in preventing attacks. To support ongoing research and facilitate community engagement, we provide open access to the study results, experimental results, and our tool, available at our website (Authors, 2024).
2. Background
We now give a basic background of notions that later sections assume readers are familiar with. In particular, a blockchain is a decentralized and distributed ledger that records transactions across multiple machines. It is constituted of a sequence of blocks that are cryptographically linked together, and each blockchain contains a set of transactions between accounts on the blockchain. A transaction refers to the act of transferring assets or information on the blockchain. In the context of Ethereum blockchain, a transaction is public and it typically involves either sending Ether (the native cryptocurrency) or an Externally Owned Accounts (EOAs) invoking or deploying a smart contract. An EOA is controlled by an individual or entity through private keys without any associated code running on the blockchain.
Smart Contract is a self-executing piece of code deployed on the blockchain that automatically enforces and executes the terms of an agreement when predefined conditions are met. Accounts that contain smart contracts are called Contract Accounts. Ethereum Virtual Machine (EVM) is the runtime environment for executing smart contracts on the Ethereum blockchain. It acts as a decentralized computing machine that processes instructions and maintains the state of all smart contracts and accounts on the network. In Ethereum, gas is a unit that measures the computational effort required to execute transactions or smart contracts. Different opcodes, have varying gas costs, with opcodes involving storage read and write being particularly expensive due to their impact on the blockchain’s state. Recently, Ethereum introduced the EVM opcodes TLOAD and TSTORE in EIP-1153 (Ethereum Improvement Proposals, 2023) that were implemented in the Cancun upgrade (Kraken Exchange, 2024) in March 2024. Similar to the standard storage read and write, TLOAD and TSTORE handle temporary data within a single transaction and reset with each new transaction. These opcodes significantly reduce gas costs for runtime validation, offering a valuable opportunity to enhance control flow integrity in smart contracts at runtime.
DeFi Protocol is a decentralized financial system consisting of interconnected smart contracts to provide services like lending, borrowing, and trading without traditional intermediaries. ERC20 is a standard template for smart contracts specifically designed for creating fungible tokens on Ethereum. The majority of valuable tokens on Ethereum are built using this standard. Etherscan is the most widely-used blockchain explorer for Ethereum. Etherscan provides labels for addresses, identifying them as hackers, MEV bots, or protocol deployers. (Eth, 2024). Control Flow in software engineering is the sequence of instructions or statements executed in a program. In the context of smart contracts, we use control flow to denote the order of function invocations within a given protocol’s contracts during a single transaction.
3. Empirical Study - Understanding Control Flows in Hacked DeFi Protocols
In this section, we present an empirical study of control flows of historical transactions of victim DeFi protocols leading up to a hack. Typically, a victim protocol operates for a period before being compromised. During this operational phase, various blockchain actors may build upon the protocol, introducing novel control flows. Our goal is to thoroughly examine these control flows and answer the following research question:
RQ1: How do control flows in hack transactions differ from those in other (benign) transactions prior to a hack?
To answer RQ1, we conducted a study on a systematically collected benchmark comprising 30 victim protocols involved in security hacking incidents. See Section 6.1 for detailed information about our benchmark collection methodology. For each hacking incident, we examine the nature and uniqueness of the control flows in the hack transactions, determining whether they differ from all previously observed transactions. The detailed study results are available in the “RQ1” column in Table 2 in Section 6.
Results: Out of the $30$ studied hack incidents, $24$ demonstrated control flows that were distinct from any previously observed transaction patterns (marked as ✓in Table 2), highlighting the novel mechanisms by which these exploits were conducted. However, in $6$ cases — specifically involving the protocols bZx, VisorFi, UmbrellaNetwork, Opyn, DODO and Bedrock_DeFi — the control flows had been observed previously in benign transactions. A detailed investigation into these exceptions revealed insightful nuances. In particular, the UmbrellaNetwork hack (marked as ✗) was traced back to an integer underflow vulnerability. The Bedrock_DeFi hack (marked as ✗) was traced back to an issue which mistakenly set the conversion ratio of ETH/uniBTC to 1:1. These two exploits were executed through one function call to steal funds, which, while not novel in terms of control flow pattern, leveraged a specific code vulnerability to breach the protocol (QuillAudits Team, 2025; UNO Re, 2021). Catching these two hacks involves combining control flow analysis with data flow analysis.
The remaining 4 hacks presented a different complexity; each involved various types of re-entrancy attacks (marked as ). These attacks, while operating under the guise of familiar top-level function calls, manifested unique and sophisticated control flows at deeper interaction levels, which are not included in the simplified top-level control flow that we extract as described in Section 6.1. This shows that even though the simple control flow-based analysis described previously can detect $24$ hacks, it misses $4$ hacks that a more refined control flow-based analysis approach may help identify. Thus, we propose to enhance our control flow analysis framework to more accurately detect and classify hacks with intricate control flows such as the above $4$ re-entrancy hacks. The specifics of this improved approach will be elaborated upon in Section 5, aiming to better capture the subtleties of complex exploit strategies and refine our understanding of hack dynamics within DeFi protocols.
Finding 1: Our findings indicate that the vast majority of hack transactions introduce unique control flows that differ from all previously observed benign transactions.
4. Preliminary and Definitions
[…]
Figure 1. Call Tree.
We now present the formalization of the control flow of a transaction which our control flow integrity technique relies on. We use protected protocol, denoted as $P$, to refer to the set of smart contracts protected by our technique. We use $C_{i}$ to represent the $i^{th}$ protected contract, where $P=\{C_{1},C_{2},\dots,C_{j}\}$. Typically, $P$ consists of all the core contracts of the protocol, as specified by the developers of the protocol. We use external contracts, denoted as $E$, to refer to smart contracts that a protected protocol is built on top of but are external to the protected protocol, such as stablecoins or oracles. In addition, any contract that is neither part of the protected protocol nor considered external is referred to as an untrusted contract (denoted as $U$).
The call tree of a transaction $tx$, denoted as $CT(tx)$, represents the tree structure of function calls made during the execution of the transaction. Each node in the tree corresponds to a function call, and each directed edge represents the call dependency between two functions. An invocation from a transaction $tx$ with respect to a protected protocol $P$, denoted as $\iota(tx,P)$, is a tuple that corresponds to a sequence of function calls starting from an entry point (a function call) associated with one of the contracts in $P$. An invocation is considered simple if it does not involve any re-entrancy to the protected protocol, i.e., between the start and end of a function call we do not have a second call to the same function. Otherwise, it is considered re-entrant. The control flow of an invocation, denoted as $CF(\iota(tx,P))$ (in short $CF(\iota)$), consists of the start and end of a function call within a protected contract, abstracting away calls between protected or external contracts that occur during the invocation. If an untrusted contract is called between two calls to protected contracts (i.e., re-entrancy to $P$), both calls to the protected contracts are included in the control flow. The control flow of a transaction $tx$ w.r.t. a protected protocol $P$, $CF(tx,P)$ is defined as a sequence of control flows of invocations. Formally, $CF(tx,P)=\langle CF(\iota_{1}),CF(\iota_{2}),...\rangle$.
[…]
5. Approach
In this section, we present the design of our control flow integrity technique. We first introduce the whitelisting policies. We then present a series of control flow simplification heuristics designed to increase the likelihood of the corresponding transaction meeting the whitelisting criteria. Lastly, we describe how these policies and heuristics are implemented within a smart contract monitor and how this monitor is applied to track transactions across different scenarios.
5.1. Control Flow Whitelisting Policies
We propose four control flow whitelisting policies to identify benign invocations.
Policy 1: Simple Independent Invocations. An invocation in a transaction is considered benign if it is simple and independent of any prior invocations executed within the transaction. The rationale for this policy is that a simple, independent invocation mirrors the behavior of a function being invoked by an EOA in a single, standalone call. This represents the fundamental usage of a function  [^1] If invoking a single function without re-entrancy within $P$ is flagged as malicious, it points to a single function access control issue, which falls outside the scope of this paper.. As such, this type of invocation is considered benign.
Policy 2: Read-Only Invocations. An invocation is considered benign if it is read-only and does not modify any blockchain state. Specifically, an invocation is read-only if it performs no write operations to the storage (i.e., no SSTORE operations), does not call other state-changing functions, and does not transfer Ether  [^2] Other operations, such as SELFDESTRUCT or CREATE, could also alter the blockchain state but are rare in high-profile DeFi protocols. If such operations are present in a function, that function should not be classified as read-only.. This is because invocations which do not alter the blockchain state have no effect on the final state. Therefore, those invocations can be omitted from the transaction without influencing the overall outcome.
Policy 3: Runtime Read-Only (RR) Function Calls. A function call is considered runtime read-only if it does not have a storage write (SSTORE), and it performs no Ether transfers. Some functions may not be marked as read-only in their source code but behave as such on the fly. An invocation is runtime read-only if all function calls within it are runtime read-only. By tracking runtime behavior and identifying such function calls and invocations, we can prune these invocations from the control flow, as they do not alter the blockchain state.
Policy 4: Restore-on-Exit (RE) Storage Writes. This heuristic permits to safely ignore storage writes that temporarily alter values but restore the original state at the end of execution, i.e., the value returned in the first read operation (SLOAD) equals the one written by its last write operation, with no write preceding the first read. This increases the likelihood of classifying function calls as runtime read-only. A typical example is the re-entrancy guard pattern, where a function restores the original state of the guard before exiting to prevent re-entrancy attacks. In our system, while these re-entrancy guards remain active and function as intended, any storage writes to them are disregarded when determining whether a function call is runtime read-only or not.
5.2. Control Flow Simplification Heuristics
To reduce the complexity of control flows and minimize the risk of flagging benign transactions as malicious, we propose two heuristics that allow CrossGuard to safely ignore certain function calls.
Heuristic 1: ERC20 Function Calls. ERC20 is a widely used smart contract standard for implementing tokens in DeFi protocols. ERC20 contracts perform token management, and several functions within these contracts modify user properties without impacting the overall protocol state. For example, the functions transfer and transferFrom change only the balances of the sender and receiver, while approve, increaseAllowance, and decreaseAllowance simply modify the allowance granted by the sender to the spender. We consider these five ERC20 functions to be benign and safe, as they have been extensively tested and are widely used across numerous DeFi projects. Therefore, calls to those functions within invocations are safely ignored.
Heuristic 2: Read-After-Write (RAW) Dependency. An invocation $\iota_{2}$ is considered storage read-after-write (RAW)-dependent on an earlier invocation $\iota_{1}$ if: $\iota_{2}$ reads from a storage location that $\iota_{1}$ writes to (with an exception of storage writes classified in Heuristic 1). In the absence of such dependencies, simple invocations are treated as independent, and a control flow consisting only of such simple and independent invocations is whitelisted. Note that re-entrant invocations, no matter it is dependent or not, will never be whitelisted by this heuristic.
5.3. System Overview
In this section, we explain how CrossGuard is integrated into a DeFi protocol pre-deployment by instrumenting the original code. The system consists of two main components: instrumentation within the protected contracts and a guard contract. Each function in the protected contracts is instrumented and assigned a unique positive integer as its function identifier.
Algorithm 1 EnterFunc and ExitFunc in guard contract
1:State Variables (accessed via tload/tsstore):
2:$CFHash,sum,invCount:int$
3:$callTrace:int[]$
4:$isCFRAW,isCFReEntrancy:bool$
5:_$\_allowedPatterns:mapping(int\to bool)$
6:function EnterFunc(funcID: int)
7:     if $sum=0$ then
8:         $invCount\leftarrow invCount+1$
9:     else
10:         $isCFReEntrancy\leftarrow true$
11:     $sum\leftarrow sum+funcID$
12:     $callTrace.push(funcID)$
13:     return $invCount$
14:function ExitFunc(funcID: int, isRR, isRAW: bool)
15:     $sum\leftarrow sum-funcID$
16:     if isRR then
17:         $callTrace.pop()$
18:     else
19:         $callTrace.push(-funcID)$
20:     if $sum=0$ then
21:         for each $id$ in $callTrace$ do
22:              $CFHash\leftarrow\text{keccak256}(id,CFHash)$
23:         $callTrace.clear()$
24:     if isRAW then
25:         $isCFRAW\leftarrow true$
26:     if $\neg$ _$\_allowedPatterns[CFHash]\land(isCFReEntrancy\lor isCFRAW)$ then
27:         revert “Unsafe pattern detected”
When an instrumented function is invoked, it sends the guard contract its identifier to record its function entry and receives a positive integer value indicating the invocation count. The instrumented function then tracks storage accesses, records storage writes using the invocation count, and sends control flow data to the guard contract upon exit. The guard contract collects this control flow data from the protected contracts, simplifies it using the defined heuristics, evaluates whether the control flow is whitelisted, and reverts the transaction if it is not.  [^3] Note when protocol administrators execute a transaction, CrossGuard includes a straightforward mechanism (not detailed but trivial to implement) that allows administrators to deactivate the lock contract at the beginning of a transaction, perform actions without interference from the control flow integrity checks, and reactivate the lock contract at the end. Algorithm 1 outlines the implementation within the guard contract, while Algorithm 2 outlines a detailed description of both the execution and instrumentation applied to the protected functions. Furthermore, Table 1 specifies the instrumentation applied to the original source code.
Algorithm in the guard contract. Algorithm 1 implements the control flow tracking in the guard contract. The function EnterFunc is invoked by protected contracts when one of their functions is called by an untrusted contract, taking a unique function identifier (funcID) as input for each function in the protected contracts. If the sum of function identifiers is zero, it signifies the start of a new invocation, and invCount is incremented (line 8). Otherwise, it indicates a re-entrancy condition, and isCFReEntrancy is set to true (line 10). The sum and callTrace are updated (lines 11-12) to record the function ID. The EnterFunc returns the invCount to the protected contracts for their future processing.
The ExitFunc (lines 14-27) 444 Both EnterFunc and ExitFunc have access control to check if the caller is a protected contract, which is omitted for brevity in Algorithm 1. is called by the protected contracts at the exit of the same function that triggered EnterFunc. It takes three arguments: funcID, isRR (isRuntimeReadOnly), and isRAW (isRead-After-Write dependent on a previous invocation). If the invocation is runtime read-only, it removes the funcID added by EnterFunc from the callTrace (line 17). Otherwise, it pushes the negated funcID onto the stack (line 19). When the sum equals zero, signaling the end of an invocation, the CFHash is computed over the entire callTrace (lines 21-22) to summarize the control flow of the invocation as a hash. Additionally, if any invocation has a RAW dependency, the algorithm sets the isCFRAW flag to true (lines 24-25). Finally, if the computed hash does not match an allowed pattern, and a read-after-write condition or a re-entrancy condition is detected, the transaction is reverted to prevent unsafe behavior, enforcing the policy to block malicious control flows (lines 26-27). Without any pre-approved control flow patterns, _allowedPatterns only include simple invocations by default. But administrators can add more patterns to this mapping to whitelist more control flows.
Algorithm 2 State Access Tracking in Protected Functions (Each step within in this algorithm is executed as part of the instrumented code, in accordance with the modifications outlined in Table 1.)
1:State Variables (accessed via tload/tsstore):
[…]
4:function ExecuteInstrumentedCode(invNum: int)
5:     $readElements,writeElements:\text{arrays of int}\leftarrow[]$
6:     $isRR:\text{bool}\leftarrow\text{true}$
7:     $isRAW:\text{bool}\leftarrow\text{false}$
8:     Execute the original source code with instrumentation outlined in Table 1.
9:     for each $slot$ in $writeElements$ do
10:         $storageWrites[invNum][slot]\leftarrow true$
11:         if $tempWrites[slot]\neq tempReads[slot]$ then
12:              $isRR\leftarrow false$
13:     for each $slot$ in $readElements$ do
14:         for $i\leftarrow 1$ to invNum do
15:              if $storageWrites[slot,i]$ then
16:                  $isRAW\leftarrow true$
17:                  break
18:     clear $tempReads$ and $tempWrites$
19:     return $isRR,isRAW$
20:function FuncUntrusted
21:     $invNum\leftarrow\textsc{EnterFunc}(\text{funcID})$
22:     $isRR,isRAW\leftarrow\textsc{ExecuteInstrumentedCode}(invNum)$
23:     ExitFunc(unique funcID, isRR, isRAW)
24:function FuncTrusted(invNum: int)
25:     $isRR,isRAW\leftarrow\textsc{ExecuteInstrumentedCode}(invNum)$
26:     return $isRR,isRAW$
Table 1. Instrumentation for Protected Functions
Original Code Instrumentation Needed
After every SLOAD ($sload(slot)\to value$): 1:$readElements.append(slot)$ 2:if $slot\not\in tempWrites$ then 3:     $tempReads[slot]\leftarrow value$
After every SSTORE ($sstore(slot,value)$): 1:$writeElements.append(slot)$ 2:$tempWrites[slot]\leftarrow value$
After every other state-changing opcode or external call to a state-changing function: 1:Set $isRR\leftarrow\textbf{false}$
After every call to other protected contracts ($funcCall()\to isSubRR,isSubRAW$): 1:$isRR\leftarrow isRR\land isSubRR$ 2:$isRAW\leftarrow isRAW\lor isSubRAW$
Algorithm for Protected Functions. Algorithm 2 implements the storage access tracking within instrumented protected functions. Table 1 provides a detailed breakdown of the instrumentation made to the original functions. Given an original function implementation, Func, it is replicated into two functions: FuncTrusted and FuncUntrusted. FuncTrusted can only be invoked by protected contracts 555 FuncTrusted has access control to check whether the msg.sender is in the whitelist of protected contracts, which is not detailed in Algorithm 2 for simplicity. , where the invNum is passed by its caller. In contrast, FuncUntrusted is designed to handle invocations from untrusted contracts or external wallets. It fetches the invNum from the guard contract by calling EnterFunc, tracks the storage accesses, and determines whether the function is runtime read-only or involves RAW (read-after-write) dependencies before sending the function ID to the guard contract via ExitFunc. Both functions call ExecuteInstrumentedCode, which tracks state changes and evaluates whether the invocation is runtime read-only and whether any read-after-write dependencies exist.
The algorithm initializes readElements and writeElements arrays to store accessed storage slots, alongside two boolean flags, isRR (runtime read-only) and isRAW (read-after-write dependencies), as described in lines 5-7 of the implementation. Then the algorithm proceeds to execute the instrumented code (line 8), with specifics provided in Table 1. During each SLOAD operation, instrumentation appends the accessed slot to readElements and records it in tempReads if it hasn’t previously been written to. Correspondingly, each SSTORE operation results in the slot being added to writeElements and its value stored in tempWrites. If any EVM opcode or function call modifies the blockchain state or if a subsequent protected contract call is not runtime read-only, isRR is set to false. Additionally, if any subsequent call to protected contracts involves a read-after-write dependency, isRAW is set to true.
Finally, the algorithm checks and updates isRR and isRAW, as well as the storageWrites variable for future invocations (lines 9-19). For each written storage slot, the slot is recorded in storageWrites (line 10). If a slot written is not restored-on-exit, the function is marked as not runtime read-only(lines 11-12). For each slot read, the algorithm checks whether the slot was written to in a previous invocation, marking the invocation as RAW-dependent if so (lines 13-17). The temporary mappings are cleared (line 18) before returning isRR and isRAW to the guard contract (line 19).
5.4. A Running Example
Figure 2 illustrates how CrossGuard integrates with a series of protected contracts to detect and prevent malicious transactions initiated by attackers. In this scenario, the attacker uses a hack contract, referred to as Contract 0 (not shown in Figure 2 for simplicity), to launch an attack on the system. The attack begins when Contract 0 invokes FuncAUntrusted in Contract 1. This action triggers EnterFunc(funcID = 1) within the guard contract, which logs the function identifier and returns invCount = 1. Since Contract 1 is a trusted contract, it proceeds to call FuncBTrusted in Contract 2, a call permissible between trusted contracts without further interaction with the guard contract. Contract 2 uses the invCount passed from Contract 1 to monitor its read-only and read-after-write dependencies. After determining that its operations are runtime read-only and lack read-after-write dependencies, Contract 1 invokes ExitFunc(1, isRR = true, isRAW = false) in the guard contract. The guard contract, recognizing the operation as runtime read-only, removes the function call from the callTrace, effectively clearing it.
Figure 2. A Running Example of CrossGuard
Subsequently, the attacker directs Contract 0 to invoke FuncCUntrusted in Contract 3, leading to a call to EnterFunc(funcID = 2) by the guard contract, which assigns invCount = 2. Contract 3, under manipulation, then triggers an interaction with Contract 4—an untrusted contract—which in turn reenters Contract 3, causing it to execute FuncDUntrusted. This interaction prompts another call to EnterFunc(funcID = 3) in the guard contract. Here, the guard contract detects potential re-entrancy (setting isCFReEntrancy to true as outlined in line 10 of Algorithm 1) and records the function identifier, returning invCount = 3. Upon completion of its operations, Contract 3 calls ExitFunc(3, isRR = false, isRAW = true). Equipped with detailed logs of the control flow and detected violations (re-entrancy and read-after-write dependencies), the guard contract opts to revert the entire transaction. This decision is taken to block the execution of the malicious operation, effectively safeguarding the blockchain network from this and similar exploits. This robust mechanism provided by CrossGuard enables real-time detection and validation of advanced threats such as re-entrancy, ensuring control flow integrity across multiple protected contracts.
5.5. Optimizations
We have implemented two optimizations to reduce gas costs.
Optimization 1: Bypassing Validation for Simple Invocations from EOAs. When a protected contract’s function does not contain a re-entrancy bug and is directly invoked by an EOA, the transaction will consistently follow a single, straightforward invocation path, which conforms to the criteria set by Policy 1 (Section 5.1). We utilized Slither (Feist et al., 2019) to identify such functions. To account for the occasional false positives reported by Slither, we also conducted a manual verification of each function. A preliminary check is implemented at the beginning of these functions. If such functions are invoked by an EOA, the control flow validation can be safely bypassed, significantly reducing gas overhead.
Optimization 2: Detecting Restore-on-Exit Storage Slots Statically. Another optimization involves statically detecting restore-on-exit storage slots. By analyzing a function’s control flow graph, we can identify certain storage slots that are restored to their original values at the end of every execution branch. If such restore-on-exit slots are detected statically, they do not need to be tracked at runtime, reducing the overhead of monitoring storage reads and writes. We implemented a prototype of this optimization on top of the open-source EVM bytecode analysis tool Heimdall (Becker, 2023). When applied to the protected contracts analyzed in Section 3, we identified $4$ contracts and $50$ functions that utilize re-entrancy guards.
6. Evaluation
In this section, we conduct a detailed assessment of CrossGuard. Our evaluation aims to answer the following research questions:
  1. RQ 2:
    How accurately does CrossGuard stop hack transactions, considering both true positives and false positives?
  2. RQ 3:
    How do various actors, aside from hackers, introduce new control flows?
  3. RQ 4:
    Can informed hackers bypass CrossGuard?
  4. RQ 5:
    What are the gas overheads of CrossGuard?
  5. RQ 6:
    How does the performance of CrossGuard compare to that of the state-of-the-art tool Trace2Inv?
Table 2. Summary of Benchmarks and Control Flow Analysis Results for Victim DeFi Protocols.
Benchmarks     RQ1     RQ3
    Unique CF in Hack?     Total #P-Tx #S-Tx #O-Tx #E-Tx
Victim Protocol Protocol Type Hack Hack Type #C         #Txs #nCF #Tx #nCF #Tx #nCF #Tx #nCF #Tx #nCF
  bZx Lending (bZx, 2025) oracle manipulation 7
29037 49 1707 24 16061 28 7736 18 3533 14
Warp Lending (War, 2025) oracle manipulation 17     ✓     416 1 11 0 404 0 0 0 1 1
CheeseBank Lending (Che, 2025) oracle manipulation 12     ✓     2404 1 125 0 1690 0 543 0 46 1
InverseFi Lending (Inv, 2025) oracle manipulation 12     ✓     126652 73 731 12 51471 17 46680 2 27770 44
CreamFi1 Lending (Cre, 2025a) re-entrancy 6     ✓     190865 121 321 32 181677 25 5795 34 3072 55
CreamFi2 Lending (Cre, 2025b) oracle manipulation 24     ✓     220962 273 500 56 204485 40 5073 33 10904 166
RariCapital1 Lending (Rar, 2025a) read-only re-entrancy 8     ✓     7142 13 58 1 6831 5 209 10 44 5
RariCapital2 Lending (Rar, 2025b) re-entrancy 12     ✓     84485 27 424 2 45217 0 27700 10 11144 19
XCarnival Yield Earning (XCa, 2025) logic flaw 6     ✓     877 3 61 0 800 0 0 0 16 3
Harvest Yield Earning (Har, 2025) oracle manipulation 11     ✓     31002 9 623 5 30338 2 2 0 39 2
ValueDeFi Yield Earning (Val, 2025) oracle manipulation 9     ✓     362 1 99 0 262 0 0 0 1 1
Yearn Yield Earning (Yea, 2024) logic flaw 7     ✓     133704 38 2855 10 81023 18 45035 8 4791 15
VisorFi Yield Earning (Vis, 2025) re-entrancy 3
86777 0 55 0 26480 0 47314 0 12928 0
UmbrellaNetwork Yield Earning (Umb, 2025) integer underflow 1     ✗     61 0 3 0 57 0 0 0 1 0
PickleFi Yield Earning (Pic, 2025) fake tokens 4     ✓     7830 2 1402 0 5168 0 1231 1 29 1
Eminence DeFi (Emi, 2025) logic flaw 6     ✓     22542 1 25 0 9473 0 9968 0 3076 1
Opyn DeFi (Opy, 2025) logic flaw 4
3937 1 29 1 607 0 1 0 3300 0
IndexFi DeFi (Ind, 2025) logic flaw 6     ✓     70735 5 98 0 14981 4 28186 0 27470 1
RevestFi Yield Earning (Rev, 2025) re-entrancy 5     ✓     2186 6 32 0 2127 5 19 0 8 1
DODO DeFi (DOD, 2025b) access control 2
1523 1 2 1 1285 0 195 0 41 0
Punk NFT (Pun, 2025) access control 3     ✓     111 4 14 1 96 2 0 0 1 1
BeanstalkFarms DAO (Bea, 2025) DAO attack 8     ✓     58648 13 243 9 43713 1 8923 0 5769 4
DoughFina Lending (Dou, 2025) no input validation 2     ✓     19 0 16 0 2 0 0 0 1 0
Bedrock_DeFi Restaking (Bed, 2025) price miscalculation 3     ✗     3426 4 3 0 2127 0 893 0 403 4
OnyxDAO DAO (Ony, 2025) fake market 8     ✓     157442 7 225 2 97210 0 393 0 59614 5
BlueberryProtocol Yield Earning (Blu, 2025) decimal difference 5     ✓     493 2 124 1 365 0 0 0 4 1
PrismaFi Restaking (Pri, 2025) no input validation 3     ✓     43669 34 135 2 24722 22 4187 0 14625 24
PikeFinance Lending (Pik, 2025) uninitialized proxy 1     ✓     8411 1 18 0 7026 0 0 0 1367 1
GFOX Game Fi (GFO, 2025) access control 2     ✓     12442 1 30 0 9392 0 12 0 3008 1
UwULend Lending (UwU, 2025) oracle manipulation 1     ✓     19765 101 178 3 18298 43 7 3 1282 84
                                  Avg. Ratio 100 100 6.81 19.7 68.34 18.4 12.67 8.92 12.18 52.78
6.1. Methodology
Hacked Protocol Selection: To address the research questions, we selected hacked DeFi protocols experiencing significant financial losses (exceeding \$300k) on Ethereum. Our dataset comprises two sets. The first set of 21 distinct victim protocols  [^6] Harvest Finance experienced two separate hacks on different contracts but is considered one protocol for our analysis. were identified by (Chen et al., 2024) from February 14, 2020, to August 1, 2022, excluding four cross-chain bridge hacks as they fall outside our scope. The second set comprises $8$ hacked protocols identified between February and July 2024 from DeFiHackLabs (Contributors, 2025). Table 2 details the selected benchmarks, presenting their Protocol Type, the hack transaction (Hack), the exploited vulnerability (Hack Type), and the number of affected contracts (#C). These vulnerabilities include diverse issues such as oracle manipulation, re-entrancy, integer underflows, access controls, and incorrect input validation. The selected protocols represent a broad spectrum of DeFi categories.
Target Contract Selection and Transaction History Retrieval: After identifying the hack transactions for each victim protocol, we collected all associated contracts involved. We used labels from Etherscan (Eth, 2024) to accurately determine the contracts belonging to the victim protocol by examining deployer addresses. We then retrieved the complete transaction history of these identified contracts from their deployment until the hack event. This comprehensive transaction history, including the hack transaction itself, constitutes the operational dataset of each victim protocol.  [^7] Although we might miss a few affected protocol contracts due to varying deployers or indirect involvement, having more contracts to protect would only introduce more complexity to the control flows, not reduce it. Thus, our analysis remains valid even with a subset of core contracts.
Control Flow Extraction and Simplification: To analyze control flows, we constructed function-level invocation trees from transaction data, aggregating invocations targeting the identified protocol contracts (detailed methodology provided in Section 4). We simplified these extracted control flows by excluding read-only functions, as they do not alter contract state and are inherently benign. Furthermore, we omitted five standardized ERC20 token functions—transfer, transferFrom, approve, increaseAllowance, and decreaseAllowance—since these functions are extensively tested, widely recognized as secure, and typically exhibit straightforward control flows without branching or external interactions. Such standard functions do not usually expose significant security risks targeted by CrossGuard. Instead, vulnerabilities emerge from more complex and interactive control flows (further discussed in Section 5.2). By applying this simplification, we concentrated on critical control flows that significantly impact protocol security, enabling a clearer understanding of the interactions leading to security breaches.
CrossGuard Evaluation(RQ2): To evaluate the effectiveness of CrossGuard, we conducted experiments under four configurations: (1) Baseline: A prototype implementing only whitelisting policies 1 and 2 (see Section 5.2). (2) Baseline+RR: Baseline augmented with Policy 3. (3) Baseline+RR+RE: Baseline augmented with Policy 3 and 4. (4) Baseline+RR+RE+ERC20: Baseline augmented with Policy 3 and 4, and Heuristic 1. (5) CrossGuard: Integrates all 4 policies and 2 heuristics into the Baseline. Note that these configurations are instrumented pre-deployment and operate autonomously post-deployment without manual intervention. CrossGuard enforces predefined policies and heuristics without relying on past transaction data. However, if an unseen control flow is mistakenly blocked, CrossGuard provides an administrative feedback mechanism that allows protocol administrators to manually approve and whitelist it. To evaluate this mechanism, we tested CrossGuard under three CrossGuard+Feedback settings, assuming administrators could approve new control flows within 3 days (19,200 blocks), 1 day (6,400 blocks), and 1 hour (267 blocks).
We assessed these configurations using historical transactions from $30$ benchmarks collected in Section 3. To further evaluate CrossGuard under extreme conditions, we applied it to another $3$ widely adopted DeFi protocols—AAVE, Lido, and Uniswap—which serve as fundamental DeFi building blocks. These protocols attract many DeFi developers and feature the most complex and continuously evolving control flows due to their high composability and extensive integrations. To conduct this evaluation, we collected the core smart contracts for these protocols from their official websites (Aave, 2024; Lido DAO, 2024; Uniswap Labs, 2024). Next, we retrieved the most recent 100,000 transactions interacting with these contracts. We then applied CrossGuard to these transactions, measuring its false positive rate (FP%) under real-world extreme conditions. The experimental results are summarized in Table 3.
DeFi Actors Identification(RQ3): To deeply understand the results of CrossGuard and the diversity of control flows in DeFi protocols, we categorize transactions according to their origins and initiators. We focus explicitly on transactions with non-trivial control flows (nCFs), as these represent complex and less predictable interactions, offering deeper insights into protocol dynamics. We identify four primary actor groups capable of introducing unique, non-trivial control flows: 1. Privileged Transactions (P-Tx): Originated by protocol deployers or administrators via privileged functions (e.g., constructors, administrative operations), typically reflecting protocol setup or administrative management activities. 2. Same Protocol (S-Tx): Transactions initiated by other contracts within the same protocol, developed internally to enhance operational coherence and overall functionality. 3. Other DeFi Protocols (O-Tx): Transactions initiated by externally deployed DeFi protocols (commonly labeled on Etherscan), often through open-source collaboration, enriching the broader DeFi ecosystem. 4. External Individuals (E-Tx): Transactions initiated by contracts deployed by external actors such as arbitrageurs, individual traders, or malicious entities (hackers), generally employing closed-source contracts to exploit vulnerabilities or execute complex strategies.
6.2. RQ2: Effectiveness of CrossGuard
Table 3. Ablation study, Gas Consumption and Bypassability of CrossGuard
RQ2 RQ5 RQ2 RQ4
Baseline +RR Baseline +RR+RE Baseline +RR+RE+ERC20 Baseline+RR+RE+ERC20 +RAW (a.k.a, CrossGuard) CrossGuard+Feedback
Baseline 3 days 1 day 1 hour
Victim Protocol Block? FPBlock? FPBlock? FPBlock? FPBlock? FPGas OH(%) Block? FPFPFPNot- Bypassable Flash- Loan
bZx2 4.51 4.51 4.5 3.88 3.57 12.38 0.5 0.31 0.12
Warp 0 0 0 0 0 0.04 0 0 0 $\ast$
CheeseBank 2.06 2.06 2.06 0 0 6.93 0 0 0 $\ast$
InverseFi 14.96 14.94 14.91 0.08 0.08 27.46 0.04 0.03 0.03 $\ast$
CreamFi1 4.30 2.28 1.16 1.15 0.39 4.93 0.13 0.07 0.03
CreamFi2 4.32 3.1 2.27 1.12 0.84 16.95 0.23 0.15 0.09 $\ast$
RariCapital1 1.92 1.25 1.25 1.23 1.22 4.91 0.64 0.53 0.41 $\ast$
RariCapital2 1.42 0.33 0.09 0.09 0.02 82.19 0.02 0.02 0.01
XCarnival 0 0 0 0 0 0.17 0 0 0
Harvest 0 0 0 0 0 0.03 0 0 0 $\ast$
ValueDeFi 0 0 0 0 0 0.07 0 0 0 $\ast$
Yearn 2.47 2.47 2.43 0.33 0 3.86 0 0 0 $\ast$
VisorFi 9.56 9.56 9.56 0 0 43.53 0 0 0
UmbrellaNetwork 0 0 0 0 0 0.07 0 0 0 N/A
PickleFi 0.71 0.71 0.71 0.01 0.01 2.08 0.01 0.01 0.01
Eminence 3.21 3.21 3.21 0 0 15.15 0 0 0 $\ast$
Opyn 0.05 0.05 0 0 0 28.74 0 0 0
IndexFi 5.67 5.64 5.64 0 0 36.85 0 0 0
RevestFi 0 0 0 0 0 36.67 0 0 0
DODO 0.07 0.07 0 0 0 0.99 0 0 0
Punk 0 0 0 0 0 0.03 0 0 0
BeanstalkFarms 5.83 5.83 5.83 0.06 0.06 30.9 0.01 0.01 0.01
DoughFina 0 0 0 0 0 0.16 0 0 0
Bedrock_DeFi 15.41 15.41 15.41 0.26 0.06 25.56 0.06 0.06 0.06 N/A
OnyxDAO 4.49 4.48 4.48 0 0 11.69 0 0 0
BlueberryProtocol 0.65 0 0 0 0 0.33 0 0 0 $\ast$
PrismaFi 31.49 31.49 31.46 1.2 0.7 7.47 0.1 0.05 0.05
PikeFinance 0 0 0 0 0 1.28 0 0 0
DFOX 3.40 3.4 3.39 0 0 30.7 0 0 0 $\ast$
UwULend 2.89 2.55 2.55 2.55 2.38 33.47 0.44 0.35 0.19 $\ast$
Summary 29 2.78 29 2.55 29 2.44 28 0.36 28 0.28 15.52 28 0.07 0.05 0.03
AAVE N/A 9.7 N/A 8.53 N/A 8.52 N/A 8.52 N/A 7.32 34.07 N/A 1.10 0.53 0.22 N/A N/A
Lido N/A 13.09 N/A 13.09 N/A 13.07 N/A 7.29 N/A 7.27 52.12 N/A 4.38 2.04 0.43 N/A N/A
Uniswap N/A 1.17 N/A 1.04 N/A 1.00 N/A 0.49 N/A 0.24 3.25 N/A 0.24 0.23 0.06 N/A N/A
Summary N/A 7.99 N/A 7.55 N/A 7.53 N/A 5.43 N/A 4.94 29.81 N/A 1.91 0.93 0.24
The columns “RQ2” in Table 3 present the results for RQ2. Each configuration is evaluated using two key metrics: “Block?” indicates whether the hack was successfully blocked; “FP%” represents the false positive rate for that configuration. Two sets of benchmarks, $30$ hacked protocols and 3 popular protocols, both include a “Summary” row at the bottom, showing the total number of blocked hacks and the average false positive rate per protocol.
Without the ERC20 heuristic, CrossGuard successfully blocks $29$ out of $30$ hacks. The only exception is UmbrellaNetwork, which falls victim to an integer underflow vulnerability (see Table 2). Since integer underflows are local bugs that do not manifest as anomalous control flows, they cannot be mitigated by control flow integrity mechanisms like CrossGuard. When the ERC20 heuristic is enabled, CrossGuard blocks $28$ out of $30$ hacks. The only exception is Bedrock_DeFi, where the attacker exploited a missing input validation vulnerability by invoking a single vulnerable function, without requiring complex control flows. In addition to this core exploit step, the attacker also executed two ERC20 functions—approve and transferFrom—which, however, were not essential to the attack itself. These ERC20 calls could have been executed as separate transactions and did not contribute to the exploit. Since removing them would not prevent the hack from occurring, CrossGuard does not block this transaction. Overall, for $28$ out of $30$ attack transactions, the exploits involve complex control flows that are effectively captured and blocked by CrossGuard, demonstrating its robustness in preventing sophisticated hacks.
Answer to RQ2: CrossGuard is highly effective, especially when combined with rapid manual feedback, resulting in significant reductions in false positives and improved system reliability.
6.3. RQ3: Control Flows Introduced by Different Actors
The columns labeled RQ3 in Table 2 summarize our analysis of non-trivial control flows introduced by each transaction category. The last row provides the average ratios of transactions and control flows for each transaction category.
A key insight from this analysis is that a significant number of protocols ($20$, as highlighted in gray) exhibit a relatively low number of non-trivial control flows ($\leq 9$) throughout their operational lifetimes prior to being hacked. Notably, in $10$ protocols (Warp, CheeseBank, ValueDeFi, Eminence, IndexFi, RevestFi, Punk, BlueberryProtocol, PikeFinance, and UwULend), the hack was the first external non-trivial control flow introduced, beyond those generated internally (as indicated by gray cells showing 0 nCF from O-Tx but exactly 1 nCF from E-Tx). This insight suggests that $11$ hacks could be simply prevented with zero false positive rates by restricting ALL external (“non-trusted”) developers, allowing only the protocol deployers to create new control flows.
The analysis reveals that E-Tx and P-Tx are significant sources of non-trivial control flows. Although external transactions account for only $12.18\%$ of the total, they introduce $52.78\%$ of non-trivial control flows, often bringing unexpected interactions and potential vulnerabilities. In contrast, privileged transactions, despite comprising $6.81\%$ of the total, contribute to $19.7\%$ of non-trivial flows, underscoring the critical role of administrative actions in protocol dynamics.
Answer to RQ3: CrossGuard effectively mitigates hacks by regulating non-trivial control flows, particularly those from external sources, demonstrating its critical role in enhancing DeFi security.
6.4. RQ4 and RQ5: Bypassability and Gas Overheads of CrossGuard
Case Studies. Given that CrossGuard operates as a fully on-chain runtime system, it is transparent, allowing attackers to study its implementations and whitelisted control flows. A prevalent concern is whether informed attackers could bypass CrossGuard by splitting complex hack transactions into simpler ones. To address this, we perform in-depth studies of the $28$ hacks blocked by CrossGuard. Our analysis involves scrutinizing the control flows, underlying vulnerabilities, and the potential outcomes if attackers were to split their transactions.
Results for Case Studies. The RQ5 column in Table 3 summarizes our findings: $25$ out of $28$ hacks cannot be bypassed by attackers. Specifically, $13$ hacks inherently require complex control flows to exploit vulnerabilities (marked as ✓).
Additionally, 12 hacks rely on executing multiple capital-intensive functions to carry out the exploit. Historically, these attacks have used flash loans, which require all steps to be completed within a single transaction. Without flash loans, the attackers have to risk significant capital while competing with arbitrage bots, a scenario we deem as non-bypassable due to the high financial risks (marked as $\ast$). Only three hacks—Punk, DoughFina, and PikeFinance (marked as ✗)—show a bypass chance for attackers. The root causes of these exploits are access control, missing input validation, and re-initialization, respectively. Each of these vulnerabilities can be exploited by invoking a single function without requiring complex control flows. These attacks were initially caught by CrossGuard because the hackers included additional preparatory operations in their transactions. However, these preparatory steps can indeed be split and executed separately in separate transactions, allowing attackers to bypass CrossGuard.
Answer to RQ4: CrossGuard’s defenses remain effective against 25 out of 28 hacks.
Experiment. To measure the gas overhead of CrossGuard, we instrumented smart contracts in a template-based manner. We insert specific code snippets at key points—such as function entry and exit, as well as during storage access operations. These snippets execute at runtime to capture the additional gas consumption introduced by CrossGuard. We used Foundry (Foundry Contributors, 2023) to compare the gas costs between the original and instrumented contracts, recording overhead differences in various execution phases. This process involved detailed assessments of each instrumentation type, functions within the guard contract, and EOA checks. During transaction replays, we logged the additional gas costs incurred at these key execution points to compute the overall gas overhead.
Results for Gas Overheads. The RQ6 column in Table 3 presents the gas overhead introduced by CrossGuard for each benchmark. On average, the overall gas overhead is $15.52\%$. Notably, $14$ protocols exhibit a gas overhead below $5\%$, primarily due to the high proportion of EOA transactions within these benchmarks. Since EOA transactions benefit from Optimization 1 (Section 5.5), which reduces unnecessary gas consumption, the resulting gas overhead remains minimal. Even for the three widely used DeFi protocols, which feature a significant number of contract-initiated transactions, the average gas overhead remains 29.81%. This is a reasonable tradeoff considering the strong security guarantees provided by CrossGuard. The higher overhead in these cases stems from the need to track function calls and storage accesses, which are essential for securing protocols with complex execution flows.
Answer to RQ5: CrossGuard demonstrates a manageable gas overhead of 15.52% on average for the hacks benchmark.
6.5. RQ6: Comparative Analysis with the SOTA Tool Trace2Inv
[…]
Table 4. Comparison of CrossGuard and Trace2Inv
CrossGuard (w/o TS) CrossGuard (w TS) Trace2Inv (w TS)
[…]
# Hacks Blocked 28 28 27 22
Avg. FP1.74 0.23 3.39 0.25
Results. Table 4 presents the analysis results that demonstrate that CrossGuard, even without training data, effectively blocks $28$ out of $30$ while maintaining an average FPrate of $1.74$%. This is a significant achievement compared to Trace2Inv, which requires training data to function but only blocks at most $27$ hacks. When trained, CrossGuard maintains its effectiveness in blocking hacks, reducing its FPrate dramatically to $0.23$%, which is superior to the FPrates achieved by Trace2Inv’s invariants ($3.39$% and $0.25$%). These results underline CrossGuard’s potential as a state-of-the-art solution providing robust security for DeFi applications. Moreover, CrossGuard and Trace2Inv can be used in conjunction to provide a more comprehensive security solution.
Answer to RQ6: CrossGuard surpasses Trace2Inv in hack prevention and false positive rates.
7. Threats to Validity
The internal threat to validity concerns potential human errors in identifying protected contracts. As discussed in Section 6.1, we rely on EtherScan labels to identify these core contracts to protect. The labels might be incomplete, leading to missing protected contracts. However, with more protected contracts identified, our approach will remain effective in blocking hack transactions, as the control flow of hack transactions will be more complex but still unique as per our approach. Our results may also face external threats due to the reliance on Trace2Inv (Chen et al., 2024) benchmarks, which focus on hacks up to June 2022. We mitigate this threat by including additional 8 hacks from February to July 2024. Additionally, we also include three major DeFi protocols—AAVE, Lido, and Uniswap—representing the most current protocols and user transactions.
8. Related Works
Invariant Generation and Enforcement. A significant body of research has focused on generating and enforcing invariants for smart contracts to ensure their security. Cider (Liu et al., 2022) employs deep reinforcement learning to derive invariants that prevent arithmetic overflows from contract source code. InvCon (Liu and Li, 2022) and InvCon+ (Liu et al., 2024) combine dynamic inference with static verification to produce verified, expressive contract invariants based on function preconditions and postconditions. Furthermore, Trace2Inv (Chen et al., 2024) dynamically learns invariants from transaction history, analyzing the effectiveness of $23$ invariants in preventing attacks. These studies predominantly focus on invariants tailored to individual contracts or specific kinds of vulnerabilities. In contrast, CrossGuard focuses on control flows across multiple contracts within the protocol. Unlike these prior studies, CrossGuard can achieve better true positive and false positive rates with a one-time configuration pre-deployment.
Control Flow Restriction. In the industry, SphereX (Spherex, 2024), a smart contract security firm, offers services to manually restrict control flows within smart contracts. While their goals align closely with ours, their approach requires protocol developers to manually select which control flows to whitelist or blacklist. This manual intervention contrasts sharply with our methodology, which not only automates the whitelisting of unseen control flows but also simplifies the overall control flow structure, significantly reducing the burden on developers and increasing the system’s adaptability.
Re-entrancy Attack Defense. Many researchers have proposed approaches to restrict smart contract control flows specifically to combat re-entrancy attacks. Tools like Oyente (Luu et al., 2016), Osiris (Torres et al., 2018), Reguard (Liu et al., 2018), Slither (Feist et al., 2019), MPro (Zhang et al., 2019), Sailfish (Bose et al., 2022), Pluto (Ma et al., 2021), Park (Zheng et al., 2022), SliSE (Wang et al., 2024), Albert et al., employ static analysis and symbolic execution to identify potential re-entrancy vulnerabilities, then apply re-entrancy guards as preventive measures. Notably, Sereum (Rodler et al., 2018) and Grossman et al. offer runtime validation frameworks to protect deployed contracts against re-entrancy attacks. Callens et al. have designed strategies to prevent the same function from being called twice within a transaction. Unlike these approaches, our work adopts a broader perspective on control flow restriction, targeting a more comprehensive set of vulnerabilities beyond just re-entrancy.
Secure Type System. Cecchetti et al. introduce a security type system that, alongside runtime mechanisms, robustly enforces secure information flow and re-entrancy controls, even amidst unknown code (Cecchetti et al., 2020, 2021). In contrast, our system is designed purely for runtime operation built entirely on the EVM, operating without the need for any supplementary type system, thereby simplifying integration and adoption.
9. Conclusion
In this paper, we presented CrossGuard, a novel control flow integrity framework specifically designed to secure smart contracts within the DeFi ecosystem. By instrumenting smart contract source code, CrossGuard dynamically prevents malicious transactions from executing risky control flow paths at runtime, effectively mitigating a broad spectrum of sophisticated attacks. Our comprehensive evaluation demonstrates that CrossGuard blocks the vast majority of benchmark attacks, significantly reduces false positives without relying on a pre-collected training set of benign transactions, and maintains a manageable gas overhead. Furthermore, integrating manual feedback enhances its accuracy, ensuring adaptability to emerging threats. Together, these results establish CrossGuard as a practical and robust solution for enhancing smart contract security, while providing valuable insights for the design of future DeFi protocols.